13.15 a. Find and CZ if z sin(xy)+ x3y b. Find evaluated at the point where u -1 and v 1, if f(x,y) x3y, x(u,v) V-u, and y(u,v) u2 +v2
Answer #1

part a:

z^2+z*x*sin(x*y)+x^3*y=0

taking derivative w.r.t. x,

2*z*(dz/dx)+x*sin(x*y)*(dz/dx)+z*sin(x*y)+z*x*cos(x*y)*y+3*x^2*y=0

==>(dz/dx)*(2*z+x*sin(x*y))=-(z*sin(x*y)+z*x*cos(x*y)*y+3*x^2*y)

==>dz/dx=-(z*sin(x*y)+z*x*cos(x*y)*y+3*x^2*y)/(2*z+x*sin(x*y))

z^2+z*x*sin(x*y)+x^3*y=0
taking derivative w.r.t. y,

2*z*(dz/dy)+(dz/dy)*x*sin(x*y)+z*x*cos(x*y)*x+x^3=0

==>dz/dy=-(z*x*cos(x*y)*x+x^3)/(2*z+x*sin(x*y))


part b:


df/du=(df/dx)*(dx/du)+(df/dy)*(dy/du)

at u=-1 and v=1

x=v-u=2

y=u^2+v^2=2


df/dx=3*x^2*y

==>df/dx=3*2^2*2=24


df/dy=x^3

==>df/dy=x^3=2^3=8

dx/du=-1

dy/du=2*u=-2

then df/du=24*(-1)+8*(-2)

=-40